# Towards Unified AI-Driven Fracture Mechanics: The Extended Deep Energy Method (XDEM)

### v1

@@ Results
This section presents the performance of the proposed Extended Deep Energy Method (XDEM) in modeling fracture evolution under varying loading and boundary conditions. The results are organized into two parts: (i) prediction of stress intensity factors (SIFs), which quantify the near-tip stress fields, and (ii) simulation of crack path propagation under different fracture modes (modes I, II and III)..
@@
Figure 2: Mixed-mode crack problem: (a) geometry, material properties, and boundary conditions; (b) comparison of XDEM-predicted stress intensity factors ($K_{I}$ and $K_{II}$) with FEM reference solutions for $a=0.5$ and $\sigma_{0}=100,\text{MPa}$ at different crack inclination angles $\beta$; (c–f) displacement and stress contours predicted by XDEM for different crack angles $\beta$ in the mixed-mode configuration.
@@
Finally, to further illustrate the generality of XDEM, Fig.˜3 presents results for intersecting cracks. XDEM accurately reproduces displacement discontinuities along multiple crack interfaces and captures stress concentration near the crack tips, in close agreement with the reference solution [16]. These results highlight the robustness of XDEM in representing complex crack topologies without requiring adaptive meshing or problem-specific refinement.
@@
It is noteworthy that, as discussed by [26], classical variational phase-field formulations are generally unable to capture crack nucleation due to the absence of an explicit representation of material strength as an independent macroscopic parameter. In contrast, XDEM naturally accommodates such effects through its energy-based formulation and extended neural representation.
@@ Discussion
The code of this work will be available at https://github.com/yizheng-wang/Research-on-Solving-Partial-Differential-Equations-of-Solid-Mechanics-Based-on-PINN after the publication of the manuscript.
@@ Acknowledgement
$\begin{cases}\sigma_{ij,j}(\boldsymbol{x})+f_{i}(\boldsymbol{x})=0&\boldsymbol{x}\in\Omega,\\ \sigma_{ij}(\boldsymbol{x})=C_{ijkl}\varepsilon_{kl}(\boldsymbol{x})&\boldsymbol{x}\in\Omega,\\ \varepsilon_{ij}(\boldsymbol{x})=\tfrac{1}{2}(u_{i,j}(\boldsymbol{x})+u_{j,i}(\boldsymbol{x}))&\boldsymbol{x}\in\Omega\setminus\Gamma^{c},\\ u_{i}^{+}(\boldsymbol{x})\not\equiv u_{i}^{-}(\boldsymbol{x})&\boldsymbol{x}\in\Gamma^{c},\\ \sigma_{ij}(\boldsymbol{x})n_{j}=\bar{t}_{i}(\boldsymbol{x})&\boldsymbol{x}\in\Gamma^{\boldsymbol{t}},\\ u_{i}(\boldsymbol{x})=\bar{u}_{i}(\boldsymbol{x})&\boldsymbol{x}\in\Gamma^{\boldsymbol{u}}.\end{cases}$ (3)
Here, $\boldsymbol{\sigma}$, $\boldsymbol{C}$, $\boldsymbol{\varepsilon}$, and $\boldsymbol{u}$ denote the stress tensor, stiffness tensor, strain tensor, and displacement vector, respectively. The domain is $\Omega$, with $\Gamma^{\boldsymbol{u}}$, $\Gamma^{\boldsymbol{t}}$, and $\Gamma^{c}$ representing the Dirichlet boundary, Neumann boundary, and crack surface, respectively. $\boldsymbol{f}$, $\bar{\boldsymbol{t}}$, and $\bar{\boldsymbol{u}}$ denote the body force, prescribed traction, and prescribed displacement. The symbol $\not\equiv$ indicates a displacement discontinuity across $\Gamma^{c}$. In this work, only Dirichlet and Neumann boundary conditions are considered, with traction-free conditions assumed on $\Gamma^{c}$. The boundaries satisfy $\Gamma^{\boldsymbol{t}}\cup\Gamma^{\boldsymbol{u}}=\Gamma$ and $\Gamma^{\boldsymbol{t}}\cap\Gamma^{\boldsymbol{u}}=\emptyset$. A schematic of the discrete crack model is shown in Fig.˜S1a.
@@
$W_{ext}=\int_{\Omega}\boldsymbol{f}\cdot\boldsymbol{u}(\boldsymbol{x})\,dV+\int_{\Gamma^{\boldsymbol{t}}}\bar{\boldsymbol{t}}\cdot\boldsymbol{u}\,dS,$
@@
XDEM in the discrete formulation faces challenges in tracking crack surfaces in three dimensions. Therefore, it is more natural and effective to employ the continuous formulation of XDEM for 3D crack problems. Here, we demonstrate the application of XDEM-C to the 3D crack configuration shown in Fig.˜S12a. The phase-field length scale is chosen as $l=0.0313$. Since the problem setup is essentially analogous to the 2D SENT test in Fig.˜S9, we use the 2D FEM phase-field solution as a reference. The load–displacement response is shown in Fig.˜S12b. It is worth noting that conventional DEM can only solve this problem if collocation points are heavily concentrated around the crack; with uniformly distributed points, DEM fails to converge. In contrast, XDEM achieves accurate results using far fewer, uniformly distributed collocation points.
@@
It should be noted that transfer can be performed incrementally between consecutive load steps. However, in the later stages of displacement-controlled loading, non-physical fracture patterns may appear (see Section˜S4.1). This issue arises due to the delicate balance between the expressive power of neural networks and the accuracy of numerical integration. Therefore, in practice, XDEM often uses the parameters obtained from the first load step as the baseline for transfer learning.

### v2

@@ Results
This section presents the performance of the proposed Extended Deep Energy Method (XDEM) in modeling fracture evolution under varying loading and boundary conditions. XDEM is presented in two forms: discrete (XDEM-D) and continuous (XDEM-C). The results discussed in the main text pertain to XDEM-D, while results for XDEM-C are provided in the Supplementary Section. The results are organized into two parts: (i) prediction of stress intensity factors (SIFs), which quantify the near-tip stress fields, and (ii) simulation of crack path propagation under different fracture modes (modes I, II and III).
@@
Figure 2: Mixed-mode crack problem: (a) geometry, material properties, and boundary conditions; (b) comparison of XDEM predicted stress intensity factors ($K_{I}$ and $K_{II}$) with FEM reference solutions for $a=0.5$ and $\sigma_{0}=100,\text{MPa}$ at different crack inclination angles $\beta$; (c–f) displacement and stress contours predicted by XDEM for different crack angles $\beta$ in the mixed mode configuration.
@@
Finally, to further illustrate the generality of XDEM, Fig.˜3 presents results for intersecting cracks. XDEM accurately reproduces displacement discontinuities along multiple crack interfaces in close agreement with the reference solution [40]. These results highlight the robustness of XDEM in representing complex crack topologies without requiring adaptive meshing or problem-specific refinement.
Figure 3: Displacement contours predicted by XDEM and FEM (reference solution) for intersecting cracks, demonstrating the accurate capture of displacement discontinuities by XDEM. The error map is computed with FEM taken as the reference solution.
@@ Discussion
The code of this work is available at https://github.com/yizheng-wang/XDEM.
@@ Acknowledgement
$w(\phi)$ Degradation function that represents the reduction of material stiffness
$g(\phi)$ Local dissipation function
$\boldsymbol{X}$ Learnable Williams series expansion in the extend function
@@
$\begin{cases}\sigma_{ij,j}(\boldsymbol{x})+f_{i}(\boldsymbol{x})=0&\boldsymbol{x}\in\Omega,\\ \sigma_{ij}(\boldsymbol{x})=C_{ijkl}\varepsilon_{kl}(\boldsymbol{x})&\boldsymbol{x}\in\Omega,\\ \varepsilon_{ij}(\boldsymbol{x})=\tfrac{1}{2}(u_{i,j}(\boldsymbol{x})+u_{j,i}(\boldsymbol{x}))&\boldsymbol{x}\in\Omega\setminus\Gamma^{c},\\ u_{i}^{+}(\boldsymbol{x})\not\equiv u_{i}^{-}(\boldsymbol{x})&\boldsymbol{x}\in\Gamma^{c},\\ \sigma_{ij}^{\pm}n_{j}^{\pm}=\bar{q}_{i}^{\pm}&\boldsymbol{x}\in\Gamma^{c},\\ \sigma_{ij}(\boldsymbol{x})n_{j}=\bar{t}_{i}(\boldsymbol{x})&\boldsymbol{x}\in\Gamma^{\boldsymbol{t}},\\ u_{i}(\boldsymbol{x})=\bar{u}_{i}(\boldsymbol{x})&\boldsymbol{x}\in\Gamma^{\boldsymbol{u}}.\end{cases}$ (3)
Here, $\boldsymbol{\sigma}$, $\boldsymbol{C}$, $\boldsymbol{\varepsilon}$, and $\boldsymbol{u}$ denote the stress tensor, stiffness tensor, strain tensor, and displacement vector, respectively. The domain is $\Omega$, with $\Gamma^{\boldsymbol{u}}$, $\Gamma^{\boldsymbol{t}}$, and $\Gamma^{c}$ representing the Dirichlet boundary, Neumann boundary, and crack surface, respectively. $\boldsymbol{f}$, $\bar{\boldsymbol{t}}$, and $\bar{\boldsymbol{u}}$ denote the body force, prescribed traction, and prescribed displacement. $\bar{\boldsymbol{q}}$ denotes the external traction vector applied on the crack surface. The superscripts $\pm$ indicate the two opposing sides of the crack. The symbol $\not\equiv$ indicates a displacement discontinuity across $\Gamma^{c}$. In this work, only Dirichlet and Neumann boundary conditions are considered, with traction-free conditions $\bar{\boldsymbol{q}}=0$ assumed on $\Gamma^{c}$. The boundaries satisfy $\Gamma^{\boldsymbol{t}}\cup\Gamma^{\boldsymbol{u}}=\Gamma$ and $\Gamma^{\boldsymbol{t}}\cap\Gamma^{\boldsymbol{u}}=\emptyset$. A schematic of the discrete crack model is shown in Fig.˜S1a.
@@
$W_{ext}=\int_{\Omega}\boldsymbol{f}\cdot\boldsymbol{u}(\boldsymbol{x})\,dV+\int_{\Gamma^{\boldsymbol{t}}}\bar{\boldsymbol{t}}\cdot\boldsymbol{u}\,dS+\int_{\Gamma^{\boldsymbol{c}}}\bar{\boldsymbol{q}}^{+}\cdot\boldsymbol{u}^{+}dS+\int_{\Gamma^{\boldsymbol{c}}}\bar{\boldsymbol{q}}^{-}\cdot\boldsymbol{u}^{-}dS,$
@@
XDEM in the discrete formulation faces challenges in tracking crack surfaces in three dimensions. Therefore, it is more natural and effective to employ the continuous formulation of XDEM for 3D crack problems. Here, we demonstrate the application of XDEM-C to the 3D crack configuration shown in Fig.˜S12a. The phase-field length scale is chosen as $l=0.0313$. Since the problem setup is essentially analogous to the 2D SENT test in Fig.˜S9, we use the 2D FEM phase-field solution as a reference. The load–displacement response is shown in Fig.˜S12b. It is worth noting that conventional DEM can only solve this problem if collocation points are heavily concentrated around the crack; with uniformly distributed points, DEM fails to converge. In contrast, XDEM achieves accurate results using far fewer, uniformly distributed collocation points. Although the accuracy of XDEM-C still shows some discrepancy compared to traditional FEM, to the best of our knowledge, the 3D load-displacement curve is the best result available in the current literature. We believe that XDEM-C still has room for improvement in the future, especially regarding the issue of overcoming the energy multiminima barrier in phase-field fracture, which is a key challenge for the future breakthroughs of XDEM-C.
@@
It should be noted that transfer can be performed incrementally between consecutive load steps. However, in the later stages of displacement-controlled loading, non-physical fracture patterns may appear (see Section˜S4.1). This issue arises due to the delicate balance between the expressive power of neural networks and the accuracy of numerical integration. Although it is fundamentally unrelated to transfer learning algorithms like LoRA, this issue primarily stems from the numerical integration precision. In the future, however, a transfer learning algorithm that improves the integration accuracy in subsequent load steps could potentially be developed for solving fracture mechanics problems using PINNs. Therefore, in practice, XDEM often uses the parameters obtained from the first load step as the baseline for transfer learning. For the fracture propagation results in the manuscript, we use the parameters obtained from the first load step as the baseline for transfer learning.